OpenStudy (anonymous):

The coordinates below represent two linear equations. How many solutions does this system of equations have? Line 1 x y –2 7 0 6 Line 2 x y 3 0 0 –3 A. 0 B. exactly 1 C. exactly 2 D. infinitely many

OpenStudy (anonymous):

i think it is d i will medal and fan and testimony who will ever help i have two more after this that are like the same plz i need to know this stuff plz explain ur answer thx :)

OpenStudy (anonymous):

@Vocaloid @calculusxy

OpenStudy (anonymous):

i meant to say i have two more after this that need checking*********

OpenStudy (anonymous):

@livias.random

OpenStudy (anonymous):

@livias.random i am tagging u

OpenStudy (anonymous):

m = (7-6)/(-2-0) m = 1/-2 m = -1/2 (y-6) = -1/2(x-0) y-6 = -1/2x y = -1/2x + 6 - m = (0-(-3))/(3-0) m = 3/3 m = 1 (y-0) = 1(x-3) y = x - 3 - x - 3 = -1/2x + 6 x + 1/2x = 6 + 3 1 1/2 x = 9 3/2 x = 9 x = 18/3 x = 6 - y = x - 3 y = 6 - 3 y = 3 - One solution: (6,3) B

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

no problem!

OpenStudy (anonymous):

:)